OpenStudy (anonymous):

solve for a . a= square root of (19-3a) +5

OpenStudy (rajee_sam):

a = √(19-3a + 5 √(19-3a) = a-5 Square both sides (19 -3a) = (a - 5)² 19 - 3a = a² - 10a +25 Combine all like terms a² - 10a +3a +25 -19 = 0 a² -7a +6 = 0 (a-6) (a-1) = 0 a = 6,1 Now put these values in your original equation and check for extraneous solutions.